Each game l0 players oot ot every 63 man inc a plover got cno eiko o out of -ine Il gama trs he played.
Answer #1

let X be the random variable denoting the number of games that a player is drug tested.

total games played by the player = 11

probability of getting drug tested = 10/53

thus, X follows a binomial distribution with parameters n = 11 and p = 10/53

to find, P(X=5)

IF X ~ Bin(n,p)

then, P(X 5) = 0.0315

so, P(X=5) = ^{11}\textrm{C}_5 \times (\frac{10}{53})^5 \times (1-\frac{10}{53})^{11-5}

or, P(X=5) = ^{11}\textrm{C}_5 \times (\frac{10}{53})^5 \times (\frac{43}{53})^{6}

P(X=5) = 0.0315